jigglypuff314 (jigglypuff314):

how would I solve for sec(arcsin(-7/25)) ?

jigglypuff314 (jigglypuff314):

so far me has sec(25/24) ?

jigglypuff314 (jigglypuff314):

@sourwing sorry for bothering you this much :/

OpenStudy (anonymous):

let y = arcsin(-7/25), the what is sin(y) ?

jigglypuff314 (jigglypuff314):

idk >,< .-.

OpenStudy (anonymous):

you just need to take sine for both sides sin(y) = -7/25

jigglypuff314 (jigglypuff314):

oh right then I drew triangle

OpenStudy (anonymous):



jigglypuff314 (jigglypuff314):

yep :) so sec(25/24) ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):



OpenStudy (anonymous):

sec(y) = 1/cos(y) = 1/(24/25) = 25/24 since y = arctan(-7/25) sec(y) = sec(arctan(-7/25)) = 25/24 <--

OpenStudy (anonymous):

you said sec(25/24) which is wrong. perhaps you meant sec(y) = 25/24 ?

jigglypuff314 (jigglypuff314):

oh right o.o so overall then sec(arcsin(-7/25)) = 25/24?

OpenStudy (anonymous):

yes

jigglypuff314 (jigglypuff314):

alright ^_^ thanks :)

OpenStudy (anonymous):

np